# Aligning Latent Spaces with Flow Priors

### v1

@@ Appendix B Additional Toy Examples
To further demonstrate the effectiveness of our proposed method, we present additional toy examples with diverse target distributions $p_{\text{data}}$: a Grid of Gaussians, Two Moons, Concentric Rings, a Spiral, and a Swiss Roll. For each of these distributions, following the visualization style of Fig. 3, we illustrate: (a) The optimized variables $\bm{y}$ (red triangles) and samples from $p_{\text{data}}$ (blue dots), overlaid on the negative log-likelihood (NLL) landscape of $p_{\text{data}}$ (background heatmap showing $-\log p_{\text{data}}(\cdot)$). (b) The landscape of the alignment loss $\mathcal{L}_{\text{align}}$ (background heatmap), with samples from $p_{\text{data}}$ (blue dots). (c) The evolution of $\mathcal{L}_{\text{align}}(\bm{y};\theta)$ (blue solid line) and the true NLL $-\log p_{\text{data}}(\bm{y})$ (red dashed line) during the optimization of $\bm{y}$.
For the Grid of Gaussians, which is also a mixture of Gaussians, the NLL $-\log p_{\text{data}}(\bm{y})$, is computed analytically. For the other distributions (Two Moons, Concentric Rings, Spiral, and Swiss Roll), where an analytical form for $p_{\text{data}}$ is not readily available, we estimate the NLL using Kernel Density Estimation (KDE). This estimation is based on $N=100,000$ samples drawn from the respective $p_{\text{data}}$ and employs a Gaussian kernel with a bandwidth of $h=0.1$. The probability density $\hat{p}_{\text{data}}(\mathbf{x})$ at a point $\mathbf{x}$ is estimated as:
@@
where $\mathbf{x}_{i}$ are the $N$ samples drawn from $p_{\text{data}}$, $d$ is the dimensionality (here, $d=2$), and $K(\cdot)$ is the Gaussian kernel function. The NLL for an optimized variable $\bm{y}$ is then approximated by $-\log\left(\hat{p}_{\text{data}}(\bm{y})\right)$. This provides an empirical measure of how well $\bm{y}$ aligns with the target distribution as estimated by KDE.
@@
Figure 5: Further illustrations of our method’s performance on various 2D toy examples. Each row corresponds to a different target distribution $p_{\text{data}}$ (Grid of Gaussians, Two Moons, Concentric Rings, Spiral, and Swiss Roll). Left column (a,d,g,j,m): Optimized variables $\bm{y}$ (red triangles) and samples from $p_{\text{data}}$ (blue dots). The background heatmap visualizes the negative log-likelihood (NLL) $-\log p_{\text{data}}(\cdot)$, with $\bm{y}$ converging to low-NLL (high-density) regions. Middle column (b,e,h,k,n): The landscape of the alignment loss $\mathcal{L}_{\text{align}}$ (heatmap) with $p_{\text{data}}$ samples (blue dots). This landscape mirrors the NLL surface, and $p_{\text{data}}$ samples are concentrated in areas of low $\mathcal{L}_{\text{align}}$. Right column (c,f,i,l,o): Training curves for $\mathcal{L}_{\text{align}}(\bm{y};\theta)$ (blue solid line) and NLL $-\log p_{\text{data}}(\bm{y})$ (red dashed line). Their strong positive correlation and concurrent decrease during optimization demonstrate that $\mathcal{L}_{\text{align}}$ effectively serves as a proxy for maximizing the log-likelihood of $\bm{y}$ under $p_{\text{data}}$.
@@ C.1 Implementation Details of the Toy Example
[…]

### v2

@@ Appendix B Additional Toy Examples
To further demonstrate the effectiveness of our proposed method, we present additional toy examples with diverse target distributions $p_{\text{data}}$: a Grid of Gaussians, Two Moons, Concentric Rings, a Spiral, and a Swiss Roll. For each of these distributions, following the visualization style of Fig. 3, we illustrate: (a) The optimized variables $\bm{y}$ (red triangles) and samples from $p_{\text{data}}$ (blue dots), overlaid on the negative log-likelihood (NLL) landscape of $p_{\text{data}}$ (background heatmap showing $\log p_{\text{data}}(\cdot)$). (b) The landscape of the alignment loss $\mathcal{L}_{\text{align}}$ (background heatmap), with samples from $p_{\text{data}}$ (blue dots). (c) The evolution of $\mathcal{L}_{\text{align}}(\bm{y};\theta)$ (blue solid line) and the true NLL $\log p_{\text{data}}(\bm{y})$ (red dashed line) during the optimization of $\bm{y}$.
For the Grid of Gaussians, which is also a mixture of Gaussians, the NLL $\log p_{\text{data}}(\bm{y})$, is computed analytically. For the other distributions (Two Moons, Concentric Rings, Spiral, and Swiss Roll), where an analytical form for $p_{\text{data}}$ is not readily available, we estimate the NLL using Kernel Density Estimation (KDE). This estimation is based on $N=100,000$ samples drawn from the respective $p_{\text{data}}$ and employs a Gaussian kernel with a bandwidth of $h=0.1$. The probability density $\hat{p}_{\text{data}}(\mathbf{x})$ at a point $\mathbf{x}$ is estimated as:
@@
where $\mathbf{x}_{i}$ are the $N$ samples drawn from $p_{\text{data}}$, $d$ is the dimensionality (here, $d=2$), and $K(\cdot)$ is the Gaussian kernel function. The NLL for an optimized variable $\bm{y}$ is then approximated by $\log\left(\hat{p}_{\text{data}}(\bm{y})\right)$. This provides an empirical measure of how well $\bm{y}$ aligns with the target distribution as estimated by KDE.
@@
Figure 5: Further illustrations of our method’s performance on various 2D toy examples. Each row corresponds to a different target distribution $p_{\text{data}}$ (Grid of Gaussians, Two Moons, Concentric Rings, Spiral, and Swiss Roll). Left column (a,d,g,j,m): Optimized variables $\bm{y}$ (red triangles) and samples from $p_{\text{data}}$ (blue dots). The background heatmap visualizes the negative log-likelihood (NLL) $\log p_{\text{data}}(\cdot)$, with $\bm{y}$ converging to low-NLL (high-density) regions. Middle column (b,e,h,k,n): The landscape of the alignment loss $\mathcal{L}_{\text{align}}$ (heatmap) with $p_{\text{data}}$ samples (blue dots). This landscape mirrors the NLL surface, and $p_{\text{data}}$ samples are concentrated in areas of low $\mathcal{L}_{\text{align}}$. Right column (c,f,i,l,o): Training curves for $\mathcal{L}_{\text{align}}(\bm{y};\theta)$ (blue solid line) and NLL $\log p_{\text{data}}(\bm{y})$ (red dashed line). Their strong positive correlation and concurrent decrease during optimization demonstrate that $\mathcal{L}_{\text{align}}$ effectively serves as a proxy for maximizing the log-likelihood of $\bm{y}$ under $p_{\text{data}}$.
@@ C.1 Implementation Details of the Toy Example
Weight Decay $\text{\times}{10}^{-4} 1.0×10$$\text{\times}{10}^{-4}$ 0.02